OpenStudy (anonymous):

what percent of 100 is24

OpenStudy (anonymous):

24%. do you understand what percent means?

OpenStudy (anonymous):

somewhat its been awhile

OpenStudy (amistre64):

24/100 = .24 , or 24% :)

OpenStudy (anonymous):

per-cent = per - century or per-100. so if I say my share of the profit is 50 percent, I mean i get 50 out of 100 or 50/100 or 0.5 of the profit.

OpenStudy (anonymous):

that is, out of every 100, my share is 50 or half of the total. if the total profit is 40, then 50 percent of 40 is half the total = 20 in other words, 50% of 40 50/100*40 = 1/2 * 40 = 20

OpenStudy (anonymous):

okay i got it and thanks

OpenStudy (anonymous):

you are welcome